OpenStudy (anonymous):

Solve (9x – 1) – (6x + 1) = 4 x = 2 x = 4 x = 1 x = 4/3

OpenStudy (anonymous):

@kentia06

OpenStudy (anonymous):

@dpgreen

OpenStudy (anonymous):

All you have to do is plug in one of the answer choices for x, and then you will find what equals 4. :)

OpenStudy (anonymous):

I GOT 2 :D

OpenStudy (anonymous):

@MiaKitty

OpenStudy (anonymous):

(9x – 1) – (6x + 1) = 4 distribute the negative then remove the parentheses and combine like terms

OpenStudy (anonymous):

It 2 @surdawi

OpenStudy (anonymous):

show me how

OpenStudy (anonymous):

9x2 = 18- 1 = 17 6x2 = 12 +1 = 13 17-13 = 4 .__.

OpenStudy (anonymous):

Ok, got it, right! :)

OpenStudy (anonymous):

Its 4 b

OpenStudy (anonymous):

No, he is right, it is 2. :)

OpenStudy (anonymous):

(9x – 1) – (6x + 1) = 4 9x-1-6x-1=4 9x-6x-2=4 9x-6x=6 3x=6 x=2

OpenStudy (anonymous):

So 2 thanks!

OpenStudy (anonymous):

yeah joe set it up the correct way

OpenStudy (anonymous):

it is 2